Ball 1, with a mass of 114 g and traveling at 11.6 m/s , collides head on with ball 2, which has a mass of 319 g and is initially at rest. Assume that ball 1 is traveling to the right before the collision and that the collision is one dimensional.(2 questions thank you)

What is the final velocity of ball 1 if the collision is perfectly elastic?

What is the final velocity of ball 2 if the collision is perfectly elastic?
Answer #1

here,

the mass of ball 1 , m1 = 114 g = 0.114 kg

initial speed of ball 1, u1 = 11.6 m/s

the mass of ball 2 , m2 = 319 g = 0.319 kg

let the final velocity of ball 1 and ball 2 be v1 and v2

using conservation of momentum

m1 * u1 = m1 * v1 + m2 * v2

0.114 * 11.6 = 0.114 * v1 + 0.319 * v2 .....(1)

and

using conservation of kinetic energy

0.5 * m1 * u1^2 = 0.5 * m1 * v1^2 + 0.5 * m2 * v2^2

0.114 * 11.6^2 = 0.114 * v1^2 + 0.319 * v2^2 .....(2)

from (1) and (2)

v1 = - 5.49 m/s

v2 = 6.11 m/s

the final velocity of ball 1 when the collison is perfectly elastic is - 5.49 m/s

the final velocity of ball 2 when the collison is perfectly elastic is 6.11 m/s